OpenStudy (anonymous):

solve equation m/-4=-2.95

OpenStudy (anonymous):

Multiply both the sides by (-4)..

Parth (parthkohli):

Try to isolate \(m\) by multiplying both sides by \(-4\).

OpenStudy (anonymous):

11.8?

Parth (parthkohli):

\(\Large {m \over -4 } \normalsize\times-4 = -2.95 \times -4 \)

Parth (parthkohli):

\( \color{Black}{\Rightarrow {\Large {m \over \cancel{-4}}} \times \cancel{-4} = m = -2.95 \times -4}\)

OpenStudy (anonymous):

Let him or her do @ParthKohli

Parth (parthkohli):

@skittles25 So, you are correct.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thank you i wasnt sure if i should of divided or multiplied

Parth (parthkohli):

You're welcome!